OpenStudy (mylahb):

Find all the solutions for the polynomials below. 4x^2+20=0

OpenStudy (cwrw238):

ok divide through by 4 x^2 + 5 = 0 x^2 = -5 there is no real number which has a negative square the roots are 2 imaginary numbers do you know what they are?

OpenStudy (mylahb):

no

OpenStudy (mathmath333):

\(\large\tt \color{black}{x^2 = -5}\) \(\large\tt \color{black}{x=\pm\sqrt{-5}=\pm\sqrt5\times i}\) \(\large\tt \color{blue}{note~~that~~\sqrt{-1}=i}\)